# Smoother-type a posteriori error estimates for finite element methods

### v2

@@ 6 Concluding Remarks
This paper develops a practical framework connecting a posteriori error estimates with smoothers of iterative linear solvers. By leveraging smoothers of available preconditioners and linear solvers, we have derived a posteriori error estimates for various PDEs such as the Poisson, curl-curl, Biot, convection-diffusion and Helmholtz equations. For symmetric problems, reliability and efficiency of smoother-type error estimators have been theoretically verified with the help of two-level stable decompositions and saturation assumptions. In addition, extensive numerical experiments demonstrate their robustness with respect to the polynomial degree.
@@ Acknowledgments
Y. L. would like to thank Prof. Shihua Gong for helpful discussion about iterative solvers for the Helmholtz equation. The authors would like to thank the anonymous referees for constructive remarks that improved the quality and presentation of this paper.

### v3

@@ 6 Concluding Remarks
This paper develops a practical framework connecting a posteriori error estimates with smoothers of iterative linear solvers. By leveraging smoothers of available preconditioners and linear solvers, we have derived a posteriori error estimates for various PDEs such as the Poisson, curl-curl, Biot, convection-diffusion and Helmholtz equations. For symmetric problems, reliability and efficiency of smoother-type error estimators have been theoretically verified with the help of two-level stable subspace decompositions and saturation assumptions. In addition, extensive numerical experiments demonstrate their robustness with respect to the polynomial degree.
@@ Acknowledgments
Y. L. would like to thank Prof. Shihua Gong for helpful discussion about iterative solvers for the Helmholtz equation. The authors would like to thank the anonymous referees for constructive remarks that improved the quality and presentation of this paper. This work was supported by the National Key R&D Program of China under grant 2024YFA1012600 and the National Natural Science Foundation of China under grant 12471346.